OpenStudy (anonymous):

Pls help I need to cancel out Y with the opposite distributive property but I don't know how to. x-3xy = 1

OpenStudy (anonymous):

At first , x-3xy=1 or, -3xy=1-x or,-y=(1-x)/3x or, y=(x-1)/3x Now , put the value of "y" in the equation and we will get --- x-3x[(x-1)/3x]=1